OpenStudy (rayman5766):

Which expression is equivalent to this one? u · 5 + u · v A. u(5 + v) B. 5(u + v) C. v(5 + u) D. u · (5 + u) · v

OpenStudy (anonymous):

a I think

OpenStudy (anonymous):

Distributive property says that: \[a \cdot b + a \cdot c = a(b+c)\] Use this property there..